Circular No. 6211

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1995X IN UGC 12160 J. Mueller reports her discovery, on a red plate taken by D. Griffith

and herself on Aug. 3 UT with the 1.2-m Oschin Schmidt Telescope in the course of the second Palomar Sky Survey, of a supernova (mag ∼ 18) located 10′′ west and 20′′ north of the center of UGC 12160 (α = 22^{h}40^{m}.8, δ = +75^{o}09′, equinox 2000.0). J. Gizis, observing with the Palomar 1.5telescope (+ CCD direct camera), measured V − I = +1.7 on Aug. 4.

A. V. Filippenko and D. C. Leonard, University of California at Berkeley, report that CCD spectra (range 310–1000 nm, resolution 0.7–1.4 nm) obtained on Aug. 24 UT with the 3-m Shane reflector at Lick Observatory show that SN 1995X is a type-II supernova. The prominent, broad Hα line has a P-Cyg profile. The narrow Hα emission line of a superposed H II region yields a redshift of 0.0038.

COMET C/1995 O1 (HALE-BOPP) Total magnitude estimates (cf. IAUC 6202): Aug. 4.62 UT, 10.5 (P.

Camilleri, Cobram, Vic., Australia, 0.20-m reflector); 14.06, 10.1 (J. E. Bortle, Stormville, NY, 0.41-m reflector); 16.95, 10.3 (J. G. de S. Aguiar, Campinas, São Paulo, Brazil, 11×80 binoculars); 18.90, 10.4 (J. Carvajal, Madrid, Spain, 0.44-m reflector); 23.47, 10.0 (Camilleri).

COMET 6P/D’ARREST Total visual magnitude and coma-diameter estimates (cf. IAUC 6197):

Aug. 1.04 UT, 8.7, 12′ (K. Sárneczky, Ráktanya, Hungary, 20×60 binoculars); 3.00, 7.9, 12′ (G. A. Milani, Eraclea, Italy, 20×80 binoculars); 18.08, 8.2, 15′ (J. G. de S. Aguiar, Campinas, Brazil, 11×80 binoculars); 19.35, 7.6, 18′ (C. S. Morris, Lockwood Valley, CA, 10×50 binoculars); 21.33, 7.9, 18′ (A. Hale, Cloudcroft, NM, 10×50 binoculars).

COMET 58P/JACKSON-NEUJMIN Total magnitude estimates: Aug. 4.93 UT, 15.0 (W. Hasubick, Buch-

loe, Germany, 0.44-m reflector); 16.56, 15.2 (A. Nakamura, Kuma, Ehime, Japan, 0.60-m reflector + CCD); 23.99, 14.5 (D. Rodriguez, Madrid, Spain, 0.26-m reflector + CCD).

1995 August 24 Daniel W. E. Green